### v2

# Jonathan Roberts1, Kai Han2, Samuel Albanie

Abstract
[…]
1 Introduction
Driven by the union of increased compute resources, a rapidly growing research community, and competition fuelled by burgeoning commercial opportunities, the capabilities of frontier large multimodal models are swiftly improving and expanding. Furthermore, the timelines for advances, especially among closed-source models, appear to be shortening, with the time between model releases diminishing. This has important implications for evaluation and benchmarking, which are critical to gauge the relative strengths and weaknesses of models. In particular, as performance increases, common benchmarks become saturated and stale – for example, GPT-4o [28] scores 90.5%, 90.2%, and 88.7% on the widely used MGSM, HumanEval, MMLU benchmarks, respectively. The usefulness of a benchmark in distinguishing model capability diminishes when there is little room for improvement. Moreover, label inaccuracies are reportedly pervasive [26] even in the most used benchmarks, further reducing possible headroom.
Figure 1: Overall performance on GRAB. Our benchmark proves challenging for frontier LMMs. The highest performing LMM attains an accuracy of just 21.7% on GRAB.
Figure 2: The GRaph Analysis Benchmark consists of 2170 synthetically generated graph analysis questions that prove challenging for frontier LMMs. The questions cover 23 graph properties organised into four core tasks: (i) Properties focuses on the analysis of features of individual functions and series; (ii) Functions and (iii) Series require computing the mean of properties across multiple functions and series; and (iv) Transforms involves determining the properties of a function after it has undergone a series of transforms.
One particular task of interest and importance is the interpretation of scientific and mathematical figures and charts, which is central to many analytical endeavours. Specifically, we target use cases where the underlying data of a figure is not accessible, for example in documents, sketches, or other image types, and quantities can only be deduced from visual interpretation. Given their prevalence, the insights gained from understanding and reasoning over these types of figures make this task an important and valuable capability for LMMs. Various forms of chart evaluation have been present in machine learning literature for quite some time through established benchmarks [15, 25] and more recent works [46, 40]. However, these benchmarks need an uplift for the current and future generations of frontier models.
To this end, we introduce GRaph Analysis Benchmark (GRAB), consisting of 2170 questions. All figures in GRAB are generated synthetically using the Matplotlib library [13]. Although some realism might be lost when solely considering synthetically generated data, there are numerous benefits to taking this approach: (1) Difficulty – the complexity of questions can be controlled and tuned; (2) Noise – ground truth is determined automatically during curation, avoiding the need for potentially erroneous post hoc annotation; (3) Composition – we can directly select the characteristics of questions in our benchmark and arbitrarily control the scale; and, (4) Contamination – while the same figure style will almost certainly have been seen during the pretraining of most LMMs, it is unlikely that the exact figures and questions generated for our benchmark have.
Leveraging the control enabled by our synthetic generation, we optimise our benchmark for usefulness and longevity. Specifically, we include challenging questions that frontier LMMs struggle to answer (see Fig. 1) and we control the aesthetic parameters of each graph to ensure all relevant features are clear and questions are answerable. We restrict the size of our benchmark to 2170 questions, balancing streamlined evaluation and robustness.
We aim for a comprehensive evaluation of the abilities of frontier models to analyse figures by including a broad range of chart types and question formats. The tasks in GRAB involve determining key data properties, such as mean, variance, and interquartile range, as well as function properties like gradients, stationary points, and function parameters. Question complexity is increased by incorporating multiple functions and series (see Fig. 2) and increasing the required answer precision. We evaluate 20 closed source LMMs on our benchmark, which proves extremely challenging for the current generation of models, with the best performing attaining a score of 21.7%. While some models exhibit a modicum of success on the easiest question set, the vast majority of our benchmark is beyond the capabilities of all models. We conduct a detailed error analysis and set of ablations to gain insights into which tasks and categories were the most challenging, as well as investigations of how question format and precision impact performance.
We summarise the contributions of this work as follows: (i) We introduce GRAB (GRaph Analysis Benchmark), a challenging 2170 question benchmark for graph analysis, (ii) We comprehensively characterise the capabilities of 20 current generation frontier LMMs on GRAB, and (iii) We provide insights into model strengths and limitations through numerous ablations.
2 Related Work
2.1 Figure Interpretation Benchmarks
Motivated by the last decade of progress in computer vision, a range of prior work has sought to evaluate figure the interpretation capabilities of leading models. One family of work has focused on synthetic data generation pipelines, profiting from the reliable ground truth and fine-grained control enabled by this approach. This has yielded figure binary classification tasks [15], captioning [7], and various structural understanding, retrieval and reasoning tasks with limited vocabularies [14]. Synthetic figure generation has also been combined with corpora of online statistics [6, 42, 25] and Kaggle datasets [46] to achieve greater figure diversity. A second line of work has sought to construct datasets by curating ‘real’ figures directly from online sources and parsing their metadata. For this purpose, figures have been sourced from research papers [41, 40] and online sources such as Our World in Data [23].
[…]
draws particular inspiration from the family of synthetic benchmarks, to which it also belongs. However, a key distinction between our tasks and prior benchmarks is GRAB’s significantly greater difficulty, as evidenced by our empirical results illustrating that even the strongest frontier LMMs struggle to make headway (Sec. 4).
2.2 LMM Evaluation and Benchmarks
Since the emergence of LMMs, numerous benchmarks have been created to evaluate and characterise their capabilities across many domains. Interpreting and answering questions of natural images is a key task, with benchmarks such as MMBench [19], SEEDBench [16], SEEDBench-2 [17], MME [9], MM-Vet [47], LVLM-eHub [45], ScienceQA [21] and MMMU [48] finding prominence. However, as outlined in [8], although widely used, some of these most popular benchmarks contain problematic questions, including those that do not require visual content to be answered. Unlike older datasets developed to cater to the supervised learning paradigm that typically required much larger scales, more recent LMM-focused benchmarks contain significantly smaller test sets of a few thousand questions or a few hundred (e.g. [35]).
Various benchmarks can be found in more niche domains, such as geographic and geospatial evaluation [39], hallucinations [12] and OCR [20]. A closely related work to GRAB is MathVista [22], which evaluates mathematical reasoning in visual contexts. MathVista is a metadataset (similar to [38]) consisting of a broad range of mostly multiple-choice questions from 31 multimodal datasets covering a diverse set of figure types and mathematical tasks ranging from table interpretation to puzzles. Though similar, the focus of MathVista on general mathematical reasoning, is distinct from ours, which is centered on the more narrow task of deriving and analysing graph properties. Moreover, our benchmark is considerably more challenging: GPT-4o scores >60% on MathVista [^1] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard compared to 13.6% on GRAB.
Statistic Number
Questions 2170
Tasks 4
Graph properties 23
Properties questions 660
Functions questions 710
Series questions 490
Transforms questions 310
10s precision questions 30
Integer precision questions 1575
1 d.p. precision questions 565
Table 1: GRAB statistics
Category Properties
Intercepts & Gradients x-intercept, y-intercept, gradient
Stationary Points stationary point(s) coordinates
Trigonometric amplitude, vertical shift, period
Functions function equations
Counting number of points, number of series
Correlation {Pearson’s, Spearman’s rank, Kendall rank} correlation coefficient
Area Bounded total area bounded, net area bounded
Measures of Spread mean, median, interquartile range, variance
Range & Extrema min/max values, domain length, range
Table 2: Categories and properties
Figure 3: GRAB category distribution
3 GRaph Analysis Benchmark
Our GRaph Analysis Benchmark consists of 2170 questions covering four core tasks and 23 graph properties. In the following section, we provide an overview of the benchmark, including descriptions of the categories and tasks, example questions and answers, and an overview of our synthetic curation process. High-level statistics can be found in Tab. 1.
3.1 Question format
All questions in our benchmark follow a single-answer format, in which each question asks for a specific property of the function(s) or data present in the graph along with required precision levels.
3.2 Categories
Questions in our GRAB benchmark cover 23 graph properties from which analytical insights can commonly be derived from data or functions in graphs and charts. We group these properties into the categories shown in Tab. 2. We consider this set of properties to cover most real use cases and require some measure of visual analytical reasoning. We refrain from including ‘easier’ OCR-centered questions – such as legend, title, axis-label interpretation – in our benchmark to keep the focus on visual analytical reasoning, and instead only include questions of this type as an ablation (see Appendix). Furthermore, as outlined in Sec. 2, numerous prior benchmarks have included this type of question.
3.3 Tasks
We structure GRAB into the following tasks covering different aspects of graph analysis. Example questions and answers for each task can be found in Fig. 4.
3.3.1 Properties
This initial task aims to provide a broad evaluation of the capability to understand and determine or derive properties of graphs, including both basic fundamental properties like number of series or more advanced properties requiring calculations, like total area bounded. Aside from number of series questions, the graphs involve just a single function or data series.
3.3.2 Functions
Functions builds on Properties with increased complexity by adding up to a total of 10 functions per graph. Questions require an additional step of averaging across all functions. For example, ‘Estimate the mean total area bounded by the function and the x-axis of the functions shown in the plot’. Added difficulty is introduced by functions overlapping.
3.3.3 Series
Similar to Functions, Series builds on the initial task by adding up to 10 data series per graph, with questions requiring estimates of the mean of specific properties across the observable series, for example, ‘Estimate the mean Pearson correlation coefficient of the data series shown in the plot’. The additional series introduces challenges such as overlapping and clustered points.
3.3.4 Transforms
Questions in the Transforms task includes a single function and require computing a graph property after performing a sequence of up to 10 transforms (rotations, translations, scales and reflections).
3.4 Curation
We leverage the Matplotlib [13] library to synthetically create the graphs used in our core benchmark and ablations studies. The following provides a brief overview of the methodology used to generate the figures and questions. For each task, we initially created a longlist of 250 questions per property. We then performed downsampling to select a final question set with uniformly distributed answers. This downselection aimed to increase the question diversity, reduced bias in the generation process and avoid an overabundance of questions with answers converging towards 0, especially for the Functions, Series and Transforms tasks.
We associate each graph property (Tab. 2) with a set of data generation functions and function parameter ranges. For categories such as Intercepts & Gradients or Functions, we sample values for the property of interest and then fit the generation parameters to produce functions/series that match the value. For the other categories, such as Correlation or Area Bounded, we randomly sample the generation parameters and then calculate the property value from the data. In each case, we ensure questions are unique. For each graph in the Properties task, aesthetic parameters (e.g. figure size, font size, and line style) are randomly sampled from a defined set of options. For the other tasks (Series, Functions and Transforms), we limit the variability of the graphs by configuring the appearance to a defined schema. As such, all graphs have the same x and y-limits, grid lines, tick labels, font and resolution (2251x2171). A single function type is used for each graph, irrespective of the number of series.
Most questions require answers to integer precision, though to increase difficulty, approximately 25% of questions require 1 decimal place precision. Questions in the Area Bounded and Functions categories are evaluated to either integer or nearest 10 precision.
Figure 4: Example questions for the four tasks in the GRAB benchmark.
3.4.1 Quality control
The curation process underwent numerous iterations of human review that sought to tune the hyperparameters involved in creating the graphs. This activity was primarily focused on jointly ensuring (i) the questions were answerable from a technical perspective, (ii) the ground-truth answers were correct and that (iii) the rendered figures were suitably readable. Similar spot-checks were carried out on the final dataset.
4 Experiments
4.1 Baselines
We evaluate 20 LMMs on GRAB, covering a broad range of both open- and closed-source models. We use chat or instruction-tuned variants rather than base models, of each model evaluated, as they are typically better at following task and output instructions. We benchmark 4 families of frontier closed-source LMMs, specifically the GPT-4 [27, 28], Gemini [10, 37] Claude 3 [3], and Reka [34] series. Additionally, we select the following open-source models for evaluation: Qwen-VL [5], TransCore-M [36], OmniLMM [32], Yi [1], CogVLM [43], and LLaVA-1.5 [18].
4.2 Experimental settings
4.2.1 Inference
We evaluate the closed-source models via the Vertex AI API [11] {Gemini, Claude}, OpenAI API [30] {GPT}, and Reka API [2] {Reka}. Following benchmarking conventions, we set model hyperparameters that result in deterministic and reproducible generation. This entails utilising a greedy search decoding strategy in which the most probable token is selected from the model vocabulary $V$ at each step, conditional on the preceding tokens i.e., $w_{n+1}=\arg\max_{w\in V}P(w|w_{1},w_{2},\ldots,w_{n})$. We achieve this by setting the $temperature$ and $topk$ parameters to zero and 1, respectively, and specifying random seeds. We run the open-source models using HuggingFace Transformers [44] and the OpenCompass toolkit [33].
4.2.2 Prompting
We adopt a simple prompting strategy throughout our evaluation, similar to that used by the popular repository simple-evals [31]. Concretely, we use a single basic user prompt consisting of the question and output format instructions. We do not modify the system prompt or tailor the prompt for each model. Although model performance is known to be sensitive to these input configurations, our investigation focuses on chat or instruction-tuned models, which are optimised for instruction following: if a model can only perform a task when a specific phrase is included in the prompt, this is a clear weakness. Being able to answer questions in a realistic setting (i.e. as a human would be presented with the questions) offers a more generalised evaluation.
4.2.3 Evaluation
Our core evaluation protocol is exact matching between model output and the ground truth answers. The only post-processing step we take is removing leading whitespace. An answer is only scored as correct if the output characters at each position match those of the expected answer. This means that although verbose answers of the form ‘The answer is…’ or ‘<Reasoning> <Answer>’ might contain the correct answer, they will be marked incorrect. While this may lower model scores compared to human or LLM evaluation (e.g. [40, 35]), we consider this a better evaluation as it jointly assesses task proficiency and instruction-following ability. Furthermore, this methodology does not necessitate (potentially) timely human scoring or the means to run a strong LLM evaluator. As mentioned previously, when considering potential downstream usage, a model is not useful if it can perform a given reasoning task perfectly but fails to reliably follow output instructions. As an ablation, we compare this strict exact matching protocol to automatic evaluation using both closed-source and open-source LLMs.
4.3 Overall results
Accuracy (%)
Task Properties Functions Series Transforms Overall
Model (660) (710) (410) (310) (2170)
Closed-source LMMs
Claude 3 Haiku [3] 14.2 6.6 8.8 3.9 9.0
Claude 3 Sonnet [3] 15.3 8.6 4.5 4.8 9.2
Claude 3.5 Sonnet [4] 41.8 15.5 11.0 10.0 21.7
GPT-4 Turbo [27] 18.5 8.5 4.9 3.5 10.0
GPT-4o mini [29] 15.8 6.8 5.7 2.9 8.7
GPT-4o [28] 24.7 10.8 9.2 3.5 13.6
Gemini 1.0 Pro Vision [10] 20.2 5.8 6.9 6.1 10.5
Gemini 1.5 Flash [37] 28.5 11.5 8.4 9.0 15.6
Gemini 1.5 Pro [37] 34.2 11.4 13.3 6.5 18.1
Reka Edge [34] 11.8 8.7 11.6 1.9 9.4
Reka Flash [34] 13.2 10.1 6.3 3.9 9.3
Reka Core [34] 1.7 0.0 4.3 0.3 1.5
Open-source LMMs
CogVLM-Chat [43] 7.0 4.9 5.1 3.9 5.4
Qwen-VL-Chat [5] 10.2 6.6 5.1 2.9 6.8
OmniLMM-3b [32] 6.7 4.9 4.1 4.5 5.2
TransCore-M [36] 7.9 9.2 7.6 3.9 7.6
Yi-VL-6b [1] 5.6 8.6 7.1 4.2 6.7
Yi-VL-34b [1] 7.6 5.9 5.5 2.3 5.8
LLaVA-1.5 7b [18] 4.7 7.5 6.5 4.8 6.0
LLaVA-1.5 13b [18] 5.0 7.7 8.4 3.9 6.5
Table 3: Overall results on the GRAB tasks. The highest open and closed-source model scores for each task are highlighted and bold.
Accuracy (%)
Category Range Measures of Spread Area Bounded Intercepts & Gradients Stationary Points Functions Correlation Extrema Counting Trigonometric
Model (360) (370) (150) (290) (160) (30) (120) (360) (60) (270)
Closed-source LMMs
Claude 3 Haiku [3] 4.7 11.9 2.7 11.4 10.6 0.0 20.8 5.8 13.3 10.0
Claude 3 Sonnet [3] 4.7 10.0 4.0 13.8 8.1 0.0 9.2 8.9 18.3 11.9
Claude 3.5 Sonnet [4] 13.6 28.6 2.7 25.5 25.6 0.0 9.2 24.7 30.0 29.3
GPT-4 Turbo [27] 6.7 13.0 0.0 13.1 11.9 0.0 5.0 8.6 18.3 14.8
GPT-4o mini [29] 3.3 10.8 2.0 10.7 8.1 0.0 2.5 8.1 26.7 15.6
GPT-4o [28] 9.2 18.6 2.0 14.1 16.2 0.0 15.8 8.1 30.0 21.5
Gemini 1.0 Pro Vision [10] 4.7 16.2 1.3 7.9 13.1 0.0 12.5 13.1 23.3 10.4
Gemini 1.5 Flash [37] 11.7 18.9 2.7 16.2 16.9 0.0 8.3 22.2 25.0 16.3
Gemini 1.5 Pro [37] 9.4 23.2 4.7 17.9 24.4 0.0 26.7 19.4 33.3 19.3
Reka Edge [34] 2.8 9.7 3.3 10.0 8.1 0.0 38.3 4.7 18.3 13.3
Reka Flash [34] 6.7 11.6 3.3 15.5 7.5 0.0 7.5 5.3 18.3 12.6
Reka Core [34] 0.0 0.0 0.0 0.0 1.2 0.0 25.0 0.3 0.0 0.0
Open-source LMMs
CogVLM-Chat [43] 6.9 5.4 2.7 4.1 5.6 0.0 0.8 4.7 16.7 7.4
Qwen-VL-Chat [5] 3.6 8.6 1.3 6.6 10.6 0.0 1.7 6.7 18.3 10.4
OmniLMM-3b [32] 8.9 3.2 2.7 5.9 3.1 0.0 0.0 5.0 11.7 6.7
TransCore-M [36] 8.6 4.1 2.7 10.0 8.8 0.0 15.8 5.8 8.3 10.4
Yi-VL-6b [1] 5.0 6.5 3.3 7.9 5.6 0.0 12.5 5.6 1.7 11.5
Yi-VL-34b [1] 2.8 7.8 2.7 8.3 1.9 0.0 8.3 4.2 6.7 10.0
LLaVA-1.5 7b [18] 9.7 4.9 3.3 4.8 5.6 0.0 9.2 5.3 5.0 6.3
LLaVA-1.5 13b [18] 8.3 4.6 2.7 7.9 4.4 0.0 18.3 5.3 8.3 5.2
Table 4: GRAB per-category performance across all tasks. The highest open and closed-source model scores for each category are highlighted and bold.
We present the overall results on GRAB in Tab. 3. A clear observation is that our benchmark proves challenging for frontier models, with low overall performances from all models (see also Fig. 1). The highest performing model, Claude 3.5 Sonnet, achieves a score of just 21.7%. As expected, the highest scores are attained on the Properties task, which includes questions that just require the derivation of a single graph property. Lower scores are achieved on the Functions and Series tasks, likely due to the added complexity of computing the mean of a given property over multiple functions/series. There is a macro-trend of slightly higher accuracy on the functions task compared to the series task, possibly due to the irregularity introduced by the noise in the data. Significantly lower accuracy is attained on the Transforms task. A contributing factor to these scores is the degree to which the models follow output format instructions. Given our exact string matching evaluation protocol, any deviation from the expected output is deemed incorrect. Despite potentially having superior mathematical reasoning abilities, models that struggle to precisely follow instructions or have proclivities to generate verbose answers suffer severe performance degradation. With the exception of Reka Core (which showed poor output format instruction following capabilities), all closed-source models outperformed the leading open-source model. However, the open-source models did consistently generate answers in the correct format.
In this low-scoring regime where even the best models struggle, a comparison of model performance should be made with the caveat that all models can barely perform the tasks in the benchmark. We anticipate our benchmark will yield most benefit when evaluating the next generation of LMMs.
4.4 Category analysis
A decomposition of model performance into question category is provided in Tab. 4. Significant variance is shown across categories, with some models much better suited for certain categories than others. Overall, the highest accuracies are attained on the Counting category, which encompasses simple questions determining the number of series shown in each plot, or the approximate number of data points. On the other hand, no model correctly answered a single question from the Functions category. This challenging category requires correct derivation of multiple parameters (e.g. the gradient and intercept for linear functions) and correct formatting of the output. Inspection of the model output suggests adherence to the strict output format was not an issue: the models were unable to correctly determine the function parameters. However, in a number of cases, proposed functions were close to the correct answers, e.g., with all parameters correct apart from one that was off by one. Minor perturbations to the graph appearance did result in a very small number of correct answers.
4.5 Evaluation protocols
Figure 5: GRAB accuracy with different automatic evaluation protocols. Performance scores are shown for each model for (i) exact matching, (ii) answer extraction with an open-source LLM then exact matching, and (iii) answer extraction with a closed-source LLM then exact matching. Accuracy differences relative to the exact matching protocol are annotated.
Figure 6: Performance on the Function, Series and Transforms tasks with increasing complexity (number of transforms or additional series/functions). Top row: accuracy across all 3 tasks with shaded region displaying 95% Wilson confidence intervals (each point on the plot represents 151 questions). Bottom row: accuracy per task decomposition.
In addition to the exact matching evaluation protocol (see Sec. 4.2.3), we explore model performance scores with more lenient automatic evaluation protocols. Specifically, we utilised an LLM to parse the generated output from the LMMs and extract the answer text, before evaluating it via exact matching as before. We carried out this evaluation method using both a frontier LLM (Gemini 1.5 Pro) and a small open-source LLM (LLaMA 3.1 8b-instruct [24]). Details of the prompts used can be found in the Appendix. Based on random sampling, the LLMs could extract specific answers from verbose output with high accuracy, however, performed poorly when provided the correct answer and prompted to actually determine the correctness of a given answer. A comparison of the different evaluation protocols can be found in Fig. 5. On the whole, there is almost no difference between the exact matching and automatic evaluation protocols. Similarly, there is little difference between extraction with a frontier closed-source and a small (8B parameter) open-source model, suggesting this approach can feasibly be conducted with little compute. However, we advocate for the adoption of exact matching without the use of an LLM as the standard evaluation approach going forwards as most models are clearly able to sufficiently follow output instructions and it provides a better signal of model capabilities. Cases where there are differences between the performance protocols indicate models that are the poorest instruction followers.
4.6 Complexity analysis
The Functions, Series and Transforms tasks were designed to include questions spanning a controllable complexity scale. For the Functions and Series tasks, complexity refers to the number of additional functions or series present on the graph, while for the Transforms task, complexity represents the number of transformations applied to the function. In both cases, the complexity score ranges from 0 (graphs with a single function or series and no transformations) to 9. In Fig. 6, we present performance decomposed along this complexity scale for a representative selection of models. Intuitively, we expect performance to decrease as complexity increases due to the added challenge of determining the quantity of interest from additional functions/series or after additional transforms. However, the complexity plots for the evaluated models follow two distinct patterns. For the better performing models, such as Claude 3.5 Sonnet and Gemini 1.5 Pro, performance does clearly decrease as complexity increases from 0 to 3, before remaining around 10%. For the weaker models, however, the results fluctuate around 10% across the entire complexity domain. In these cases, even the lowest complexity questions are too challenging. As expected, the questions requiring decimal precision proved much harder than integer precision questions.
4.7 Question format
We carry out an ablation analysing the effect of question format on model performance by re-evaluating the questions from the Properties task in a multiple-choice setting with 5 candidate answers. The results of this comparison are displayed in Tab. 5. Multiple-choice options were generated adversarially by sampling values close to the correct answer.
For the majority of models, accuracy scores are higher in the multiple-choice setting, and most models score above the random chance score. However, the highest attained score is only 32.3%, further reflecting the difficulty of the GRAB benchmark. For a few models, including the two leading models, the opposite is true. In these cases, it is possible that the presence of plausible incorrect answers (i.e., close to the true answer) can cause confusion and lead the models to make incorrect selections.
Accuracy (%)
Model Single-answer Multiple-choice
Random chance - - 20.0
Closed-source LMMs
Claude 3 Haiku [3] 14.2 +9.4 23.6
Claude 3 Sonnet [3] 15.3 +10.5 25.8
Claude 3.5 Sonnet [4] 41.8 -10.0 31.8
GPT-4 Turbo [27] 18.5 +11.3 29.8
GPT-4o mini [29] 15.8 +12.2 28.0
GPT-4o [28] 24.7 +3.0 27.7
Gemini 1.0 Pro Vision [10] 20.2 -7.5 12.7
Gemini 1.5 Flash [37] 28.5 +2.4 30.9
Gemini 1.5 Pro [37] 34.2 -1.9 32.3
Reka Edge [34] 11.8 +11.5 23.3
Reka Flash [34] 13.2 +7.6 20.8
Reka Core [34] 1.7 +5.6 7.3
Open-source LMMs
CogVLM-Chat [43] 7.0 +14.1 21.1
Qwen-VL-Chat [5] 10.2 +11.3 21.5
OmniLMM-3b [32] 6.7 +13.6 20.3
TransCore-M [36] 7.9 +12.9 20.8
Yi-VL-6b [1] 5.6 -1.5 4.1
Yi-VL-34b [1] 7.6 +5.6 13.2
LLaVA-1.5 7b [18] 4.7 +14.2 18.9
LLaVA-1.5 13b [18] 5.0 +17.4 22.4
Table 5: Accuracy on the Properties task with different question formats. Score differences between the formats are shown as coloured text. The highest open and closed-source model scores for each format are highlighted and bold.
5 Conclusions
We introduce GRAB, a GRaph Analysis Benchmark for LMMs. We leverage synthetic data generation to create high-quality, difficult questions that prove challenging for frontier LMMs, with the best-performing model attaining a score of 21.7%. On GRAB, we evaluate a suite of 20 closed and open-source models, attaining a broad profile of the abilities of the current generation of LMMs. During inference, we prompt the models using lightweight prompts with clear output format instructions and assess answer correctness by comparing the output string with the ground truth. This approach jointly assesses the abilities of models to perform the GRAB reasoning tasks and reliably follow output instructions, enabling straightforward evaluation without the need for human or automatic evaluation using LLMs. A consequence of this harsh evaluation protocol is that some models commonly regarded as ‘frontier’ – that are prone to verbose generation – perform relatively poorly. We directly compare performance results using exact matching evaluation with automatic LLM evaluation and find little difference, suggesting most current models are able to sufficiently follow output instructions, and therefore validating exact matching as a suitable evaluation protocol. Additionally, we conduct a series of studies revealing the categories the models struggle with and the impact of different question types. We release our dataset and evaluation codebase for the community to use and hope our work encourages research towards the next generations of LMMs.
Acknowledgements
This work was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (reference EP/S022961/1), an Isaac Newton Trust grant, a research gift from Google, an EPSRC HPC grant, the Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27208022), and HKU Seed Fund for Basic Research. Samuel would like to acknowledge the support of Z. Novak and N. Novak in enabling his contribution.
Appendix
We structure this Appendix into 4 parts. In (1), we include examples of the prompts used for inference and LLM evaluation. Then, in (2) we include details of the compute costs of our work. Next, in (3) we present results on an OCR-based graph analysis question set. Finally, in (4) we detail the specific versions of the API models used in this work, aiding reproducibility.
Appendix A Prompts
Inference prompt:
[…]
LLM evaluation prompt:
[…]
Appendix B Compute
As we restricted the size of our benchmark to just 2170, the total compute required for this work is relatively low. Negligible compute was needed to create the synthetic images. Inference with the closed-source models was carried out via API calls. Using a single NVIDIA A100-80GB GPU, inference on the entire GRAB benchmark using LLaVA-1.5 7b can be carried out in approximately 30 minutes, using a single process (inference time is significantly reduced with multiprocessing).
Appendix C OCR experiments
We construct a small set of 35 OCR questions based on text in the legend, title, and axis labels. The graphs are created using a similar pipeline to the Properties task except rather than using purely randomly selected data generation processes, we use plausible data ranges and functions for the axis label. We leverage GPT-4 Turbo to construct a database of ‘realistic’ dependent and independent variables along with their ranges, distributions and directions (e.g. linear, increasing). We evaluate these questions in a multiple-choice setting with 5 options and present the results in Tab 6. Compared to performance on the Properties task with multiple-choice options (Tab. 5), the models attain much higher scores on these OCR-style questions, with many models achieving either 100% or close to 100% accuracy. These high scores suggest this particular question type is too easy for current frontier LMMs, therefore we refrain from including it in GRAB.
Model Accuracy (%)
Random chance 20.0
Closed-source LMMs
Claude 3 Haiku 42.9
Claude 3 Sonnet 91.4
Gemini 1.0 Pro Vision 100.0
Gemini 1.5 Flash 100.0
Gemini 1.5 Pro 100.0
GPT-4 Turbo 97.1
GPT-4V 97.1
GPT-4o 82.9
Reka Core 97.1
Reka Edge 88.6
Reka Flash 100.0
Open-source LMMs
TransCore-M 80.0
Yi-VL-6b 51.4
OmniLMM-3b 65.7
Qwen-VL-Chat 71.4
Table 6: Accuracy scores on OCR experiments.
Appendix D API model versions
These are the specific versions of the API models used in this work:
  • GPT-4 Turbo:
    gpt-4-turbo-2024-04-09
  • GPT-4o mini:
    gpt-4o-mini-2024-07-18
  • GPT-4o:
    gpt-4o-2024-05-13
  • Gemini Pro Vision:
    gemini-1.0-pro-vision-001
  • Gemini 1.5 Flash:
    gemini-1.5-flash-preview-0514
  • Gemini 1.5 Pro:
    gemini-1.5-pro-preview-0514
  • Claude 3 Haiku:
    claude-3-haiku@20240307
  • Claude 3 Sonnet:
    claude-3-sonnet@20240229
  • Claude 3.5 Sonnet:
    claude-3-5-sonnet@20240620
  • Reka Edge:
    reka-edge-20240208
  • Reka Flash:
    reka-flash-20240226
  • Reka Core:
    reka-core-20240501
  • Gemini-Pro:
    gemini-1.0-pro-001